(w) the average of the 39 u,s . national parks under 900,000 acress (in thousand of acres) is shown here . construct a frequency distribution for the data using eight classes.

41 66 233 775 169
36 338 233 236 64
183 61 13 308 77
520 77 27 217 5
650 462 106 52 52
505 94 75 265 402
196 70 132 28 220
760 143 46 539
1
Expert's answer
2020-01-16T12:02:20-0500

Minimum value in data set = 5.

Maximum value in data set = 775.

Range = 775 - 5 = 770.

Number of classes = 8.

Width of class interval = 770/8 = 96.25 which can be rounded up to 97.

Ordered data:

5

13

27

28

36

41

46

52

52

61

64

66

70

75

77

77

94

106

132

143

169

183

196

217

220

233

233

236

265

308

338

402

462

505

520

539

650

760

775

Now we need to count number of data in appropriate interval like 17 data are between class interval 5 - 101.

Relative frequency is calculated by dividing frequency of class to total frequency.

Like for class 5 - 101, the relative frequency = 17/39 = 0.44.

Frequency Distribution: